The price elasticity for rice is estimated to be -0.4 and the income elasticity is 0.8. At a price of $0.40 per pound and a per capita income of $20,000, the demand for rice is 50 million tons per year.


 (a) Is rice an inferior good, a necessity, or a luary? Explain 


(b) If per capita income increases to $20,500, what will be the quantity demanded of rice?


(c) If the price of rice increases to $0.41 per pound and income per capita remains $20,000, what will be the quantity demanded?


1
Expert's answer
2021-02-09T07:03:34-0500

(a) Since the income elasticity is positive therefore it can be concluded that the good is not an inferior good. Furthermore,the price elasticity of demand and income elasticity are both inelastic indicating that the good must be a necessity.

(b) IEoD = (% Change in Quantity Demanded)/(% Change in Income)

0.8=(% Change in Quantity Demanded)/0.025

(% Change in Quantity Demanded) "=0.025 \times 0.8"

=0.02

=2 %

New quantity = "50 \times (1+2 \%)"

New quantity = 51 million tons

(% Change in Income)= "\frac{20500-20000}{20000}"

"=2.5 \%"

(c) PEoD = (% Change in Quantity Demanded)/(% Change in Price)

-0.4=(% Change in Quantity Demanded)/0.025

% Change in Quantity Demanded="0.025 \times (-0.4)"

% Change in Quantity Demanded= -0.01 = -1%

New quantity= "50 \times (1-1 \%)"

New quantity= 49.5 Million Tons

(% Change in Price)= "\frac{0.41-0.4}{0.4}=0.025"